Determine the molarity M of a solution of 0.150 g of NaOH in 100 mL of water (Na = 23; O = 16; H = 1)
1
Expert's answer
2020-11-20T10:03:50-0500

CM=n/V

n=m/M

M (NaOH)=40 g/mol

n (NaOH)=0.150/40=0.004 mol

100 ml=0.1 L

CM(NaOH)=0.004/0.1=0.04 M